# Optimization of Superconducting Niobium Nitride Thin Films via High-Power Impulse Magnetron Sputtering

### v1

@@ V Conclusions
In reactive DC sputtered NbN films, it has been shown that a hexagonal crystal phase with a lower $T_{\mathrm{C}}$ begins to form as one approaches a $1{:}1$ stoichiometric ratio.Benkahoul et al. (2004); Wright, Xing, and Jena (2023) We have found that, for reactive HiPIMS, the nitrogen gas concentration that produces the highest $T_{\mathrm{C}}$ is greater than that for DC sputtering, and the maximum $T_{\mathrm{C}}$ for the HiPIMS films is greater than the maximum $T_{\mathrm{C}}$ for the DC sputtered films. This holds true across all film thicknesses studied, from $6$ nm to $120$ nm. We believe that this is the result of the HiPIMS process enabling the films to get closer to optimal stoichiometry before beginning to form the hexagonal $\epsilon$-NbN crystal phase. The lower normal-state resistivity of the HiPIMS films is attributed to increased adatom mobility, and hence improved film crystallinity, enabled by the higher kinetic energy in the HiPIMS process.
@@
The HiPIMS deposition process presents a large parameter space, and further exploration of this space may be productive. For example, we have not yet studied the effect of varying parameters such as the deposition chamber pressure and the height and duration of the HiPIMS voltage pulses. In future work, we also plan to explore the benefits of the HiPIMS process for other materials in the superconducting transition metal nitride family.
@@ Acknowledgements.
This work was supported by the National Science Foundation through awards ECCS-2117007 and ECCS-2000778.

### v2

@@ V Conclusions
In reactive DC sputtered NbN films, it has been shown that a hexagonal crystal phase with a lower $T_{\mathrm{C}}$ begins to form as one approaches a $1{:}1$ stoichiometric ratio.Benkahoul et al. (2004); Wright, Xing, and Jena (2023) We have found that, for reactive HiPIMS, the nitrogen gas concentration that produces the highest $T_{\mathrm{C}}$ is greater than that for DC sputtering, and the maximum $T_{\mathrm{C}}$ for the HiPIMS films is greater than the maximum $T_{\mathrm{C}}$ for the DC sputtered films. This holds true across all film thicknesses studied, from $6$ nm to $120$ nm. We believe that this is the result of the HiPIMS process enabling the films to get closer to optimal stoichiometry before beginning to form the hexagonal $\epsilon$-NbN crystal phase. The lower normal-state resistivity of the HiPIMS films is attributed to improved film crystallinity enabled by the higher kinetic energy in the HiPIMS process. These observations are consistent with XRD analysis, which show that the HiPIMS films have a larger lattice constant, consistent with greater nitrogen concentration, and a narrower peak width, consistent with a larger crystallite size.
@@
The HiPIMS deposition process presents a large parameter space, and further exploration of this space may be productive. For example, we have not yet studied the effect of varying parameters such as the deposition chamber pressure and the height and duration of the HiPIMS voltage pulses, including the positive kick pulse. In future work, we also plan to explore the benefits of the HiPIMS process for other materials in the superconducting transition metal nitride family.
@@ Acknowledgements.
This work was supported by the National Science Foundation through awards ECCS-2117007 and ECCS-2000778. We thank Drs. T.J. Mullen and G.A. Wurtz for performing AFM characterization.